OpenStudy (loser66):

How to find cross product of 2 vectors in R^4 ?

OpenStudy (loser66):

\(\left[\begin{matrix} 1\\1\\0\\0\end{matrix}\right]\) and \(\left[\begin{matrix} -1\\0\\1\\0\end{matrix}\right]\)

OpenStudy (loser66):

Let v is the required vector, so that v orthogonal to both vector above. It is easy to see that \(v= \left[\begin{matrix} 0\\0\\0\\1\end{matrix}\right]\) satisfies the condition. But I don't want it. I want the method to find it out, not just trial and error. Please, help me out

OpenStudy (amistre64):

as far as i can recollect, cross product is only defined in R^3 but I do recall a paper ina journal that discussed the usefulness of a nxm determinant where n /= m

OpenStudy (loser66):

So, how can we find it out in this case?

OpenStudy (loser66):

we can use dot product also. right? Let v =<a, b, c, d> , hence v . v1 = < a, b, c, d> .< 1,1,0,0> = 0 (we need it) and v. v2 = < a, b, c, d> .< -1,0,1,0> =0 also. But it is hard to find out the result

OpenStudy (amistre64):

yes, dotproduct will determine if a vector is parralel to another.

OpenStudy (loser66):

for the first one, it is a + b =0 the second one , it is -a + c = 0 hence a = -b, and c, d are arbitrary, a = c and b, d are arbitrary then v can be < 1, -1, 1, 0> , but it doesn't satisfy v.v1 =0 or v.v2 =0 :(

OpenStudy (amistre64):

the plane in R^4 is most likely teh sae approach as a circle to a sphere, add on the missing dimension

OpenStudy (loser66):

I know my logic is flaw at somewhere, but don't know how to fix yet

OpenStudy (amistre64):

n = (a,b,c,d) a(x-xo) + b(y-yo) + c(z-zo) + d(ko) = 0 but then im thinking that in R^4 there are multiple 'normals' to a plane

OpenStudy (amistre64):

k-ko that is

OpenStudy (loser66):

oh, i got it. the first one gives me <a, -a, 1, 1> since I can pick any of c, d the second one gives me <a, 1, a , 1> since i can pick any of b and d But they must be the same. that is < a, -a, 1, 1> =< a, 1, a, 1> hence a = a, -a =1 , 1 = a. it is possible if and only if a =0, hence the required v is v =< 0,0,0,1> Am I right?

OpenStudy (amistre64):

abcd 1100 ----- a+b = 0 abcd -1010 ------ -a+c= 0 a = a b = -a c = a d = d

OpenStudy (loser66):